OpenStudy (anonymous):

If (9x − 4)(9x + 4) = ax 2 − b, what is the value of a?

OpenStudy (anonymous):

@ganeshie8 @p@prowrestler @dan815 @sammixboo @Jhannybean @mathmath333

Vocaloid (vocaloid):

foil method (9x-4)*(9x+4) = 81x^2-16

OpenStudy (anonymous):

a can be 18 because 18x can =ax*2-b if b=1

OpenStudy (anonymous):

Okay, so @DaBest21 says 18. What about you @Vocaloid?

Vocaloid (vocaloid):



Vocaloid (vocaloid):

find a

OpenStudy (anonymous):

6561...?

Vocaloid (vocaloid):

 look at the underlined parts

OpenStudy (anonymous):

A=81?

Vocaloid (vocaloid):

right

OpenStudy (anonymous):

Okay, cool. Thanks! I will let you know if it is correct.